OpenStudy (anonymous):

Amanda exercised for 10 minutes every day in the first week, 20 minutes in the second week, 30 minutes in the third week, and 40 minutes in the fourth week. Billy exercised for 5 minutes every day in the first week, 10 minutes in the second week, 20 minutes in the third week, and 40 minutes in the fourth week.

OpenStudy (anonymous):

Which statement best describes the methods used by Amanda and Billy to increase the time they spent exercising? Amanda's method is linear because the number of minutes increased by an equal number every week. Billy's method is linear because the number of minutes increased by an equal factor every week. Both Billy and Amanda's method is exponential because the number of minutes increased by an equal factor every week. Both Billy and Amanda's method is exponential because the number of minutes increased by an equal number every week.

OpenStudy (anonymous):

thanks can you help me with one more plz

OpenStudy (anonymous):

Oops, accidentally removed the answer. Sure

OpenStudy (anonymous):

A store had 100 t-shirts. Each month, 30% of the t-shirts were sold and 25 new t-shirts arrived in shipments. Which recursive function best represents the number of t-shirts in the store? f(n) = f(n - 1) • 0.3 + 25, n > 0 f(n) = 100 - f(n - 1) • 0.3 + 25, n > 0 f(n) = f(n - 1) • 0.7 + 25, n > 0 f(n) = 100 - f(n - 1) • 0.7 + 25, n > 0

OpenStudy (anonymous):

nevermind i got the answer

OpenStudy (desiraem):

what's the answer to the first question?